2 2 X Questions 17 and 18 will deal with the linear system X = - 23 pts Question 17 What are the eigenvalues to the linear system? Select the correct answer O 0,-4 O 2,2 O-2 2i 0,4 O 22i4 pts Question 18 The solution of the linear system is Select the correct answer O None of the above (C X ce2 sin(2t) e cos(2
Answer #1

2 Let A= 2 -2 2 The characterisHe polynomial of A is 2 AlA=Aae A)At dat A =4A8 A(A) =0 =→ ペー4ス+8 0 - 4±4と 6-32 2. Henae igenv18 Let V= emvectofA Coenpendig jo = 2-2s an Then (A- AIV= O o) 2 22 2 2. ba 2ia 2b 0 - Let a Ken b2 envectosA (sorenpondirg f-2it 1 e -22t CD2t)-2t) 2-22t 2t 2t Now V llol-inet 24)-n2t).] 2 t 2t et) i e at n(2t) o2t(t) and X2 |(4m)asp êneanindependen